the ratio of boys to girls in a class is 2:3 if there are 18 girls in the class, how many boy are there
Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

the ratio of boys to girls in a class is 2:3 if there are 18 girls in the class, how many boy are there
-----
2:3 is the same as 2x:3x where 3x is the number of girls.
----
3x = 18
x = 6
--------
2x = # of boys = 2*6 = 12
==============
Cheers,
Stan H.

Answer by checkley79(3341)

2/3=x/18 cross multiply.
3x=2*18
3x=36
x=36/3
x=12 the number of boys.
Prof:
2/3=12/18 reduce right hand fraction.
2/3=2/3